### v1

# FedSA-GCL: A Semi-Asynchronous Federated Graph Learning Framework with Personalized Aggregation and Cluster-Aware Broadcasting

Abstract
Federated Graph Learning (FGL) is a distributed learning paradigm that enables collaborative training over large-scale subgraphs located on multiple local systems. However, most existing FGL approaches rely on synchronous communication, which leads to inefficiencies and is often impractical in real-world deployments. Meanwhile, current asynchronous federated learning (AFL) methods are primarily designed for conventional tasks such as image classification and natural language processing, without accounting for the unique topological properties of graph data. Directly applying these methods to graph learning can possibly result in semantic drift and representational inconsistency in the global model. To address these challenges, we propose FedSA-GCL, a semi-asynchronous federated framework that leverages both inter-client label distribution divergence and graph topological characteristics through a novel ClusterCast mechanism for efficient training. We evaluate FedSA-GCL on multiple real-world graph datasets using the Louvain and Metis split algorithm, and compare it against 9 baselines. Extensive experiments demonstrate that our method achieves strong robustness and outstanding efficiency, outperforming the baselines by an average of 2.92% with the Louvain and by 3.4% with the Metis.
keywords:
Federated graph learning , Asynchronous federated learning , Non-IID data , Client clustering
journal: Expert Systems with Applications
[…]
{highlights}
Address efficiency and heterogeneity in semi-asynchronous Federated Graph Learning.
Propose Federated Semi-Asynchronous Graph Cluster Learning (FedSA-GCL).
Structurally clustered clients benefit from peers despite no upload.
Achieves +3.4% accuracy and +3.1% robustness over baseline.
Achieves 484.5 fewer trips over baseline, improving communication efficiency.
1 Introduction
Graphs are data structure for modeling complex relational systems across a wide range of domains, including social networks(Guo & Wang, 2020), drug discovery(Liu et al., 2022)(Wang et al., 2023), and recommendation platforms(He et al., 2023)(Yu et al., 2022). Although graph machine learning has made significant progress in recent years, most existing methods rely on centralized storage of large-scale graph data on a single server. However, with growing concerns about data security and user privacy, this centralized model is often impractical in real-world scenarios. In practice, graph data is typically distributed across multiple data owners, and cross-site data collection becomes impractical due to privacy restrictions(Fu et al., 2022).
To address this challenge, the concept of Federated Graph Learning (FGL) has emerged, which integrates federated learning (FL) with graph learning. FGL enables collaborative model training among distributed graph data holders without requiring raw data exchange, thereby preserving privacy. In FGL, graph data is typically partitioned into local subgraphs held by different clients, such as transportation subnet works (Zhang et al., 2021b; Li & Zhu, 2021), social communities (Guo & Wang, 2020), or financial chains (Yang et al., 2021). A key challenge for FGL is subgraph heterogeneity, which refers to the significant divergence in label distributions and graph structures—driven by strong node dependencies—across clients(Baek et al., 2023; Xie et al., 2021), thereby undermining cross-client model generalization. To mitigate the impact of subgraph heterogeneity, recent studies in FGL have begun to explicitly account for the unique structural and distributional characteristics of graph data. For example, several works incorporate topology-aware designs to capture local graph semantics (Zhang et al., 2021a; Li et al., 2024b), while others explore personalized modeling or representation-level adaptation to better handle client-specific variations (Tan et al., 2022; Baek et al., 2023; Li et al., 2024a).
Despite their effectiveness, these methods are still primarily designed under synchronous training protocols, which pose significant bottlenecks to efficiency in practical settings (Jiang et al., 2022). Consequently, current FGL research continues to face two major limitations.
L1: Synchronous FGL Suffers from Inefficiency. Although existing FGL methods have demonstrated promising results (Baek et al., 2023; Li et al., 2024b, a; Zhu et al., 2024), they are predominantly designed for idealized synchronous settings. However, their reliance on synchronization raises serious concerns for practical deployment. In heterogeneous environments, such as those involving edge devices or clients with limited computing capacity, stragglers can delay the entire training process, leading to prolonged communication rounds and underutilized system resources (Chen et al., 2019). Additionally, in real-world scenarios, client participation is often sporadic and inconsistent. Strictly requiring all selected clients to complete training before aggregation is incompatible with such dynamic participation patterns, rendering synchronous protocols not only inefficient but often infeasible.
L2: Asynchronous Settings Amplify Graph Non-IID. Existing asynchronous federated learning (AFL) methods (Xie et al., 2019; Xu et al., 2023; Ma et al., 2021; Nguyen et al., 2022) are primarily developed for domains such as computer vision and natural language processing. These approaches typically assume independent and identically distributed (IID) data, and lack dedicated mechanisms to address the cross-node dependencies and topological heterogeneity intrinsic to graph data. Although SWESALT (Liao et al., 2023) introduces asynchronous communication into FGL, it does not explicitly address the Non-IID nature of graph-structured data. The inherent non-IID nature of graph data arises from topological heterogeneity and cross-node dependencies. While synchronous FL can mitigate some of this heterogeneity by aggregating updates from all clients simultaneously, asynchronous participation disrupts this balance. The server often aggregates updates from a temporally sparse and biased subset of clients, causing certain structural patterns to be underrepresented. This amplifies the existing non-IID challenge, further degrading the consistency and generalizability of the global model.
To address the above issues, we propose Federated Semi-Asynchronous Graph Cluster Learning (FedSA-GCL), a novel and scalable optimization strategy that performs effective client clustering based on their node label distribution characteristics. For similar clients, we assign aggregation weights by jointly considering their topological characteristics and model staleness. Furthermore, leveraging the asynchronous nature of the system, we introduce a ClusterCast mechanism that proactively pushes model updates to inactive clients sharing similar graph structures.
In summary, the main contributions of this paper are: (1) New Paradigm. To the best of our knowledge, this is the first paper to investigate the federated graph learning under asynchronous settings, targeting the limitation 1 and limitation 2 where existing Synchronous FGL approaches. (2) Novel Method. We propose FedSA-GCL, a semi-asynchronous federated learning method for graphs that improves cross-client consistency through a cluster-level broadcasting mechanism that mitigates stale updates by enabling non-uploaded but structurally similar clients to benefit from peer updates. (3) State-of-the-art Performance. Extensive experiments on 7 benchmark graph datasets demonstrate the superiority of FedSA-GCL compared with FedGTA baselines interms of effectiveness (3.4% accuracy improvement), robustness (3.1% accuracy improvement),and communication efficiency(484.5 client trips improvement).
The paper is organized as follows. In Section 2, we present the preliminaries and related work, including background knowledge on federated learning and a review of recent advances in federated graph learning and asynchronous federated learning. Section 3 introduces the proposed FedSA-GCL framework, detailing its core components and mechanisms. Section 4 describes the experimental setup and provides a comprehensive analysis and discussion of the results. Finally, Section 5 concludes the paper with a summary of findings.
2 Preliminary and Related Work
2.1 Problem Formulation
Due to the lack of publicly available benchmark datasets for federated graph learning (FGL), existing works (Zhang et al., 2021a; Chen et al., 2024; Baek et al., 2023) typically simulate FGL scenarios by splitting a single global graph into multiple subgraphs. Given a global graph $G=(\mathcal{V},\mathcal{E})$ with $|\mathcal{V}|=n$ nodes and $|\mathcal{E}|=m$ edges, its structure can be represented by an adjacency matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$. Each node $v\in\mathcal{V}$ is associated with a feature vector $\mathbf{x}_{v}\in\mathbb{R}^{f}$, forming the feature matrix $\mathbf{X}\in\mathbb{R}^{n\times f}$, and a one-hot label vector $\mathbf{y}_{v}\in\mathbb{R}^{|\mathcal{Y}|}$, resulting in the label matrix $\mathbf{Y}\in\mathbb{R}^{n\times|\mathcal{Y}|}$, where $|\mathcal{Y}|$ is the number of classes. To simulate a federated environment, community detection algorithms such as Metis(Karypis & Kumar, 1998) or Louvain(Blondel et al., 2008) are applied to split the global graph $G$ into $N$ subgraphs. Each client $i\in\left\{1,\dots,N\right\}$ is assigned a local subgraph $G_{i}=(\mathcal{V}_{i},\mathcal{E}_{i})$, along with the corresponding local adjacency matrix $\mathbf{A}_{i}\in\mathbb{R}^{|\mathcal{V}_{i}|\times|\mathcal{V}_{i}|}$, feature matrix $\mathbf{X}_{i}\in\mathbb{R}^{|\mathcal{V}_{i}|\times f}$, and label matrix $\mathbf{Y}_{i}\in\mathbb{R}^{|\mathcal{V}_{i}|\times|\mathcal{Y}|}$.
2.2 Federated Graph Learning
FedAvg (McMahan et al., 2017) is the most widely used optimization strategy in federated learning. It performs global model aggregation by computing a simple weighted average of the model parameters received from participating clients, where the weights are typically proportional to the size of local training data. In this setting, we integrate graph neural networks (GNNs)(Kipf & Welling, 2016) with FedAvg. The aggregation process at round $t$ is given by Eq. 1.
$\tilde{\omega}^{t}=\sum_{i=1}^{N}\frac{n_{i}}{n}\cdot\omega_{i}^{t-1},$ (1)
where $\tilde{\omega}^{t}$ is the global model at round $t$, $\omega_{i}^{t-1}$ is the local model from client $i$ in the previous round, $n_{i}$ is the number of training nodes on client $i$, and $n=\sum_{i=1}^{N}n_{i}$ is the total number of training nodes.The local update on client $i$ during round $t$ is defined in Eq. 2.
$\omega_{i}^{t}=\tilde{\omega}^{t}-\eta\cdot\nabla f(\tilde{\omega}^{t},G_{i}(\mathcal{V}_{i},\mathcal{E}_{i})),$ (2)
where $\tilde{\omega}^{t}$ is the received global model, $\eta$ is the learning rate, and $\nabla f(\cdot)$ denotes the gradient of the local loss function over the subgraph $G_{i}$.
Although FedAvg is widely adopted, it performs poorly under non-IID client data and fails to capture the unique properties of graph-structured data. To address this, FedSage+(Zhang et al., 2021a) integrates GraphSAGE into the federated setting and reconstructs missing inter-subgraph edges, but mainly focuses on structural completion. FedProto(Tan et al., 2022) tackles label distribution heterogeneity via prototype sharing, yet ignores graph topology. FedPub (Baek et al., 2023) introduces functional embeddings and sparse aggregation to enable personalized training on heterogeneous subgraphs. FedGTA (Li et al., 2024b) and AdaFGL (Li et al., 2024a) incorporate topological confidence and structure-aware personalization strategies, improving both convergence and adaptability. More recently, FedTAD (Zhu et al., 2024) adopts topology-aware, data-free distillation to enable structure-sensitive aggregation without data sharing.
Despite these advancements, most federated graph learning (FGL) methods rely on synchronous communication, where global aggregation is blocked until all selected clients complete local training. This limits scalability and leads to inefficiencies in heterogeneous or unstable environments.
2.3 Asynchronous Federated Learning
To overcome the inefficiencies of synchronous federated learning under system heterogeneity, Asynchronous Federated Learning (AFL) has been proposed to enable non-blocking client-server communication (Xu et al., 2023). AFL can be broadly categorized based on its aggregation strategy into fully asynchronous and semi-asynchronous approaches. Fully asynchronous methods such as FedAsync (Xie et al., 2019) and TWAFL (Chen et al., 2019) allow the server to update the global model immediately upon receiving a single client update, improving reactivity but risking instability under non-IID conditions. In contrast, semi-asynchronous methods like FedSA (Ma et al., 2021) and FedBuff (Nguyen et al., 2022) trigger aggregation only after collecting a certain number of updates, striking a balance between convergence stability and communication efficiency. TimelyFL (Zhang et al., 2023) introduces a tier-based scheduling mechanism that adapts local training based on device availability and latency, further enhancing responsiveness in heterogeneous environments.
Although some preliminary efforts, such as SWESALT (Liao et al., 2023), have extended semi-asynchronous strategies to graph data, they do not incorporate graph-specific optimizations and overlook the non-IID nature of graph data. Current AFL methods typically assume IID input and lack mechanisms to address graph-specific challenges such as topological heterogeneity and cross-node dependencies. As a result, their direct application to graph learning often leads to unstable convergence and limited generalization, highlighting the need for graph-aware asynchronous federated algorithms.
3 FedSA-GCL Framework
Figure 1: Overview of the FedSA-GCL framework.
FedSA-GCL is a semi-asynchronous federated learning framework specifically designed for graph-structured data, where the server initiates aggregation once a predefined number of client updates are received. As illustrated in Fig. 1, the framework follows a classic client-server architecture, enhanced with several novel designs to address non-iid data and partial communication. Clients are first clustered based on their soft label feature matrices (SFM), and model aggregation is performed by incorporating local smoothness confidence (LSC) and model staleness, promoting better representation alignment across clients. Crucially, our ClusterCast mechanism enables structurally similar but inactive clients to benefit from peer updates, effectively mitigating the degradation commonly seen in fully asynchronous settings. (1) The left panel illustrates the client-side workflow, including the generation of SFM, the calculation of LSC, and the handling of asynchronous model updates received during local training via the ClusterCast mechanism. (2) The right panel presents the server-side workflow, including the similarity-based client clustering process and the aggregation of client models weighted by LSC and model staleness.
In the following sections, we elaborate on each core design: Section 3.1 introduces the construction of SFM and the SFM-based clustering algorithm. Section 3.2 details the construction of LSC and the staleness-aware aggregation strategy; Section 3.3 explains how the ClusterCast mechanism leverages client clustering to support semi-asynchronous propagation.
3.1 Client Clustering Based on Soft Label Feature Matrix
Compared with hard labels, soft labels can better capture uncertainty and inter-class similarity in categorization. Therefore, when modeling client-side label features, we adopt node soft labels as the basis.The soft label of nodes is computed by passing input features through an encoder followed by softmax activation, as defined in Eq. (3).
$\hat{Y}=Softmax(f(\tilde{\omega_{i}};G_{i}(\mathcal{V}_{i},\mathcal{E}_{i});))$ (3)
where $G_{i}(\mathcal{V}_{i},\mathcal{E}_{i})$ denotes the local subgraph of client i, and $f(\tilde{\omega_{i}};\cdot)$is the predictive model parameterized by the aggregated weights $\tilde{\omega}_{i}$.
Since connected nodes exhibit similarity in both feature distributions and labels, while different clients present distinct label distribution patterns, we construct a Soft label Feature Matrix for adjacent nodes by multiplying their soft label features based on each client’s graph data. Typically, nodes closer to the client center (which generally have higher degrees) better represent the client’s label distribution. Therefore, we weigh the adjacent nodes’ soft label feature matrix using node degrees as weights, thereby enhancing the contribution of critical nodes. Finally, the client’s overall soft label feature representation is obtained by aggregating all weighted soft label feature matrices. The Soft label Feature Matrix(SFM) is computed as defined in Eq. (4):
$SFM=\sum_{(i,j)\in\epsilon}\left(D_{i}\times D_{j}\right)\left\langle\hat{Y}_{i},\hat{Y}_{j}^{T}\right\rangle$ (4)
where $\epsilon$ denotes the set of all edges in this client, $i$ and $j$ are the two nodes connected by an edge, $D_{i}$ and $D_{j}$ represent their degrees, and $\hat{Y}_{i}$ and $\hat{Y}_{j}$ are their respective soft label vectors.
After obtaining the SFMs from all clients, the server queries all SFMs stored in the Server Knowledge Base and computes the pairwise similarity between clients. The Server Knowledge Base maintains client history information uploaded to the server, which contains the latest client model parameters and other information. For client pairs whose similarity exceeds the predefined threshold $\theta$, their model parameters undergo mutual aggregation, and the target clients for parameter distribution are subsequently determined. Based on SFMs, the similarity between two clients is computed using cos similarity, as shown in Eq. (5). The specific calculation process is shown in Fig. 1, Feature Clustering Module.
$\begin{gathered}\mathcal{I}_{i}=\{j\mid\operatorname{sim}(i,j)\geq\theta\}\cup\{i\},\quad\forall i,j\in\operatorname{Set}(N),\ j\neq i\\ \operatorname{sim}(i,j)=\frac{\sum_{p=1}^{k*K}\operatorname{SFM}_{i}^{p}\cdot\operatorname{SFM}_{j}^{p}}{\sqrt{\sum_{p=1}^{k*K}\left(\operatorname{SFM}_{i}^{p}\right)^{2}}\cdot\sqrt{\sum_{p=1}^{k*K}\left(\operatorname{SFM}_{j}^{p}\right)^{2}}}\end{gathered}$ (5)
where $\text{SFM}_{i}^{p}$ denotes the $p$-th element in the flattened soft label matrix of client $i$, $\mathcal{I}_{i}$ is the set of clients similar to client $i$, and $\theta$ is the similarity threshold.
3.2 Model Aggregation Based on Local Smoothness Confidence and Staleness
To quantify the topological-attribute correlation between nodes, we introduce k-step Non-parametric Label Propagation (Non-param LP), which models the relationship between target nodes and their k-hop neighbors through iterative updates(Gasteiger et al., 2018).To capture higher-order structural information, the $k$-step label propagation is given in Eq. (6).
$\hat{Y}_{i}^{(k)}=\lambda\cdot\hat{Y}_{i}^{(0)}+(1-\lambda)\cdot\sum_{j\in\mathcal{N}(i)}\frac{1}{\sqrt{D_{i}D_{j}}}\cdot\hat{Y}_{j}^{(k-1)}$ (6)
where $\hat{Y}_{i}^{(0)}$ is the initial soft label of node $i$, $\mathcal{N}(i)$ denotes the 1-hop neighbors of node $i$, $D_{i}$ is the degree of node $i$, and $\lambda\in[0,1]$ controls the propagation balance.
The entropy of node soft labels quantifies the prediction consistency of local models, while the local smoothing confidence characterizes their consistency with neighborhood predictions(Li et al., 2024b). Specifically: Lower entropy indicates stronger prediction consistency within local graph structures, yielding more reliable learned topological features. Furthermore, considering the significant influence of high-degree nodes on local smoothing confidence, we assign them higher weights. The local smoothness confidence (LSC), which reflects the entropy-based consistency between nodes and their neighbors, as shown in Eq. (7).
$LSC$ $=\sum_{i=1}^{|\mathcal{V}|}D_{i}\cdot\left(e^{-1}-Entropy(\hat{Y}_{i}^{(k)})\right)$ (7)
$=\sum_{i=1}^{|\mathcal{V}|}D_{i}\cdot\left(e^{-1}+\left(\sum_{j=1}^{|\mathcal{Y}|}\hat{Y}_{i,j}^{(k)}\cdot\log\hat{Y}_{i,j}^{(k)}\right)\right)$
where $|\mathcal{V}|$ is the number of nodes, $|\mathcal{Y}|$ is the soft label dimensionality (number of classes), and $D_{i}$ is the degree of node $i$.
In asynchronous federated learning frameworks, clients conduct local training based on different versions of the global model, resulting in model version staleness(Xie et al., 2019). To address this issue, we employ a staleness metric to quantify version discrepancy and accordingly adjust client model aggregation weights. The staleness-aware aggregation weight is given in Eq. (8):
$\tilde{\omega}=\sum_{i=1}^{n}\frac{LSC_{i}\cdot(t-\tau_{i}-1)^{-\alpha}}{\sum\nolimits_{j=1}^{n}\cdot\left(LSC_{j}\cdot(t-\tau_{j}-1)^{-\alpha}\right)}\cdot\omega_{i}$ (8)
where $t$ is the current server round, $\tau_{i}$ is the last upload round of client $i$, $\omega_{i}$ is its model, and $\alpha$ is a staleness attenuation coefficient.
3.3 Semi-Asynchronous ClusterCast Broadcasting
A key challenge in semi-asynchronous federated learning is maintaining a timely and coherent information flow under partial client participation. Existing methods like FedBuff(Nguyen et al., 2022) trigger aggregation upon receiving a fixed number of client updates, but only send the aggregated model to those clients who participated in that current round. As illustrated in Fig. LABEL:fig:comm_diagram_sub1, this limitation can cause structurally similar but inactive clients—such as client 3 in round 2—to miss out on valuable updates, ultimately leading to slower convergence and increased inter-client drift. To address this issue, we propose ClusterCast, a novel broadcasting mechanism designed to enhance cluster-level cross-client consistency and training efficiency.
ClusterCast proactively pushes model updates to both uploaded clients and inactive ones that share similar graph structures. As shown in Fig. LABEL:fig:comm_diagram_sub2, this allows structurally similar clients, such as client 3, to benefit from timely updates even without participating in the current round. This design offers two key advantages:(1) It enhances information sharing among semantically similar clients;(2) It accelerates convergence.
Specifically, during each communication round, the server aggregates a personalized model $\omega_{i}$ for each uploaded client $i$ based on its semantic cluster $\mathcal{I}_{i}$ (Eq.(5)) and optimized model parameters (Eq.(8)). The resulting $\omega_{i}$ is then sent not only to client $i$, but also to all $j\in\mathcal{I}_{i}\setminus\mathcal{U}$, where $\mathcal{U}$ denotes the current uploaded set. This allows semantically similar clients to benefit from peer updates even without participating in the current round. To ensure training continuity, clients receiving updates during local training do not overwrite their models immediately. Instead, they cache the incoming model and, upon training completion, perform a confidence-aware aggregation with their local model, as defined in Eq. (9).
$\omega_{\text{i}}^{\prime}=\frac{\tilde{LSC}}{\tilde{LSC}+LSC_{i}}\cdot\tilde{\omega}+\frac{LSC_{i}}{\tilde{LSC}+LSC_{i}}\cdot\omega_{i},\tilde{LSC}=\sum_{j=1}^{|\mathcal{I}_{i}|}LSC_{j}$ (9)
where $\tilde{\omega}$ is the downloaded server model, $\omega_{i}$ is the current model at client i, $|\mathcal{I}_{i}|$ is the cardinality of $\mathcal{I}_{i}$ , $\tilde{LSC}$ and $LSC_{i}$ are the aggregated and local LSC values, respectively.
Algorithm 1 FedSA-GCL: Server Procedure
1:Initialize: $t\leftarrow 0$, $\mathcal{U}\leftarrow\emptyset$, $\mathcal{B}\leftarrow\emptyset$triangleright Global round, uploaded client set, broadcast set
[…]
4:     $t\leftarrow t+1$,$\mathcal{U}\leftarrow\emptyset$$\mathcal{B}\leftarrow\emptyset$
[…]
6:         Receive $(LSC_{i},SFM_{i},\omega_{i},\tau_{i},i)$ from UploadQueue.get()
7:         Update ServerKnowledgeBase with $(LSC_{i},SFM_{i},\omega_{i},\tau_{i})$
8:         $\mathcal{U}\leftarrow\mathcal{U}\cup\{i\}$
9:     end while
10:     for each client $i\in\mathcal{U}$ do $\triangleright$ Distribute personalized aggregated model to each uploaded client
11:         Compute aggregation set $\mathcal{I}_{i}$ via Eq. (5)
12:         Compute aggregated model $\omega_{i}$ via Eq. (8)
13:         Send $(\omega_{i},t)$ to $\text{DownloadQueue}_{i}$
14:         $\tilde{LSC}_{s}\leftarrow\sum\nolimits_{c\in\mathcal{I}_{i}}LSC_{c}$,  $\omega_{s}\leftarrow\omega_{i}$
15:         $\mathcal{B}\leftarrow\mathcal{B}\cup(\mathcal{I}_{i}\setminus\mathcal{U})$
16:     end for
17:     for each client $s\in\mathcal{B}$ do $\triangleright$ Broadcast aggregated model to semantically similar non-uploaded clients
18:         Send $(\omega_{s},t,\tilde{LSC}_{s})$ to $\text{DownloadQueue}_{s}$
19:     end for
20:end while
Algorithm 2 FedSA-GCL: Client Procedure
[…]
3:     Receive $(\tilde{\omega},\tau,\tilde{LSC})$ from DownloadQueue.get()
4:     Replace stored update with $(\tilde{\omega},\tau,\tilde{LSC})$triangleright Keep only the latest update
5:end while
6:if $\tilde{LSC}=\texttt{null}$ then
7:     $\omega\leftarrow\tilde{\omega}$
8:else
9:     Compute $\omega$ via Eq. (9)
10:end if
11:Train model $\omega^{\prime}$ using local data and $\omega$
12:Compute soft label via Eq. (3)
13:Compute $SFM$ via Eq. (4)
14:Compute non-parametric LP via Eq. (6)
15:Compute $LSC$ via Eq. (7)
16:Send $(\omega^{\prime},\tau,SFM,LSC,\text{client ID})$ to UploadQueue
The ClusterCast mechanism is seamlessly integrated into the FedSA-GCL protocol through coordinated client and server routines. The algorithm flow of FedSA-GCL on the server side and client side is shown in Algorithm 1 and Algorithm 2.
To ensure clarity and robustness under the semi-asynchronous setting, we highlight several key conventions: (1) Each client maintains a thread-safe download queue and retains only the most recent update, discarding stale ones before training; (2) Clients apply server updates only after completing local training to avoid conflicts between overlapping computation phases; (3) Each server update is timestamped with $\tau$ to guide the computation of local smoothness confidence (LSC) and aggregation weighting; (4) For each uploaded client, structurally similar but non-uploaded clients are grouped into a broadcast set and receive proactive model pushes; (5) Local aggregation is selectively performed only when the server update includes structural confidence information ($\tilde{LSC}$), enabling consistency-aware adaptation.
On the server side (Algorithm 1), lines 13–15 identify non-uploaded but semantically similar clients and proactively broadcast the aggregated models to them. On the client side (Algorithm 2), lines 2–5 process all incoming updates and retain only the latest one for local training. These coordinated routines enable timely model propagation and alignment within structurally coherent client groups.
ClusterCast is the core mechanism that enables FedSA-GCL to operate effectively under semi-asynchronous conditions, ensuring that model alignment is maintained across similar clients, regardless of their upload frequency. It significantly mitigates stale update problems and accelerates convergence, particularly in heterogeneous federated graph scenarios.
4 Experiments
In this section, we conduct extensive experiments to validate the effectiveness of our proposed method. We first introduce seven graph datasets used for training, describe the experimental setup in detail, and split the data into 20 clients using the Metis and Louvain community detection algorithms. Federated graph learning is then performed on nine baseline methods. Subsequently, we aim to answer the following research questions: Q1:In asynchronous scenarios, can our method achieve better predictive performance than state-of-the-art baselines on non-IID data? Q2:Where does the performance improvement of our method come from? Q3:How robust is our method in asynchronous federated graph learning settings? Q4:Does our method demonstrate a higher computational efficiency under asynchronous?
4.1 Experiments Setup
Dataset. To comprehensively evaluate the performance of our method against other baselines, we select seven graph datasets. Our experiments are conducted on three small-scale citation networks (Cora, CiteSeer, PubMed) (Yang et al., 2016), two medium-scale user-item datasets (Amazon Computers, Amazon Photo), and two medium-scale co-authorship networks (Coauthor CS, Coauthor Physics) (Shchur et al., 2018). Detailed descriptions of these datasets and the proportions of the train/validation/test splits are presented in Table 1.
To simulate realistic federated graph learning scenarios, we apply two community detection algorithms, Metis(Karypis & Kumar, 1998) and Louvain(Blondel et al., 2008), to split each dataset into 20 clients. These splits are used for federated graph training. For several representative datasets and both community detection algorithms, the label distributions across clients are illustrated in Fig. 3.
Table 1: The statistical information of the experimental datasets.
dataset nodes features edges classes train/val/test description
Cora 2708 1433 5429 7 20%/40%/40% citation network
CiteSeer 3327 3703 4732 6 20%/40%/40% citation network
PubMed 19717 500 44338 3 20%/40%/40% citation network
Amazon Photo 7487 745 119043 8 20%/40%/40% co-purchase graph
Amazon Computer 13381 767 245778 10 20%/40%/40% co-purchase graph
Coauthor CS 18333 6805 81894 15 20%/40%/40% co-authorship graph
Coauthor Physics 34493 8415 247962 5 20%/40%/40% co-authorship graph
Figure 3: Client-wise label distributions for selected datasets split via Metis and Louvain, demonstrating the resulting non-IID data heterogeneity.
Baselines. We compare our method against nine baselines. For traditional federated learning (FL), we select FedAvg(McMahan et al., 2017), FedProx(Li et al., 2020), and MOON(Li et al., 2021). For federated graph learning (FGL), we include FedSage+(Zhang et al., 2021a), FedGTA(Li et al., 2024b), and FedTAD(Zhu et al., 2024). For asynchronous federated learning (AFL), we compare with FedAnsy(Xie et al., 2019), TWAFL(Chen et al., 2019), and FedBuff(Nguyen et al., 2022). All experiments are conducted using the GCN model as the backbone architecture.
Hyperparameters. Our asynchronous training scripts are developed based on the open-source federated graph learning framework OpenFGL(Li et al., 2024c). For baseline methods, we adopt the hyperparameter settings reported in their original papers. In federated learning, the number of communication rounds to reach a target accuracy is a common evaluation metric(Nguyen et al., 2022). However, since asynchronous and synchronous approaches differ in their notion of a round, FedBuff proposes the client trip — a cycle of model download, local training for one epoch, and model upload—as a fair basis for comparison. Following this practice, we also adopt the client trip as the evaluation unit to ensure fair comparison across methods. We set the number of local epochs to 1 and train for 2000 client trip. To simulate an asynchronous environment, 30% of clients are randomly designated as edge devices, which fall behind the global server by 2 to 5 complete training cycles (e.g., with 20 clients, one global training cycle corresponds to 20 client trip). For cluster-based federated learning, the clustering threshold is set to 0.5, as suggested in FedGTA. Model validation is performed locally on each client’s local dataset.
Experiment Environment. All experiments are conducted on a machine equipped with an Intel Core i9-13900K processor @5.8GHz and an NVIDIA GeForce RTX 3090 GPU with 24GB memory, running CUDA 12.1. The operating system is Ubuntu 22.04, and the system has 64GB of RAM.
4.2 Performance Comparison
To answer Q1, we conduct experiments on various datasets split into 20 clients using different community detection algorithms. Each baseline is trained five times, and we report the mean performance along with the 95% confidence interval. Our method outperforms the baselines in the vast majority of cases. Specifically, under the Louvain splitting, our method achieves an average improvement of 2.92% over the baselines, and under the Metis splitting, the average improvement is 3.4%. Detailed results are presented in Table 2 and Table 3, where the best results are highlighted using bold.
Table 2: Performance comparison under Louvain splitting.
type Method Cora CiteSeer PubMed CS Photo Computers Physics Overall
FedAvg $69.0_{\pm 1.3}$ 71.8±0.5 82.9±0.6 79.1±3.2 67.2±5.3 62.1±7.2 88.0±1.3 73.9±3.1
FL FedProx 67.6±1.4 71.2±0.3 82.8±0.7 78.3±1.5 64.1±3.8 62.9±2.0 88.9±2.1 73.2±3.2
MOON 56.4±1.3 69.2±5.8 65.6±1.2 70.4±1.7 66.5±5.4 64.0±3.5 85.8±1.2 68.3±3.0
FedSage+ 68.3±1.7 69.6±0.6 81.7±4.3 78.7±3.2 72.9±6.2 65.6±3.8 88.5±0.5 75.1±2.8
FGL FedGTA 74.0±2.7 71.4±0.5 82.6±0.4 88.4±0.2 82.4±0.6 73.9±0.6 93.7±0.2 80.2±2.6
FedTAD 63.6±9.8 70.9±0.4 77.2±8.2 71.0±1.3 65.8±5.5 62.7±2.9 87.0±0.3 71.2±3.1
FedAsyn 73.4±0.7 70.9±0.4 71.1±0.4 74.7±1.5 77.5±0.9 70.1±0.7 88.3±0.4 75.1±2.1
AFL TWAFL 69.3±0.7 47.8±1.0 64.1±1.2 61.5±3.6 74.0±1.8 66.0±2.1 77.5±1.7 65.7±3.2
FedBuff 73.8±1.0 70.6±0.3 84.1±0.1 81.2±0.9 73.3±4.4 69.3±2.2 92.0±1.4 77.8±2.7
FedSA-GCL 75.8±1.6 71.8±0.5 85.0±0.2 85.8±0.1 87.6±1.5 82.6±1.0 92.9±0.5 83.1±2.3
Table 3: Performance comparison under Metis splitting.
type Method Cora CiteSeer PubMed CS Photo Computers Physics Overall
FedAvg 63.0±2.7 71.6±0.9 81.8±0.4 71.0±3.0 74.3±6.9 56.6±5.2 87.9±0.3 72.4±3.8
FL FedProx 56.7±4.4 71.5±0.8 82.3±0.6 70.3±2.8 73.7±5.3 57.6±6.2 87.8±0.4 71.4±4.1
MOON 70.4±4.3 71.1±0.5 77.9±6.5 85.0±2.2 74.8±0.8 66.1±2.5 87.8±0.5 76.1±2.7
FedSage+ 73.7±0.8 69.6±0.7 82.5±1.8 78.5±0.9 80.9±0.2 71.6±1.7 89.4±1.3 78.0±2.3
FGL FedGTA 73.2±2.4 71.4±0.7 79.5±2.3 89.3±0.2 82.9±1.3 73.2±0.8 94.0±0.3 81.0±2.9
FedTAD 50.3±8.5 71.2±0.7 70.2±6.5 67.7±3.2 70.2±5.9 63.6±3.8 86.9±1.0 68.6±3.7
FedAsyn 77.7±0.3 71.2±0.2 76.6±1.1 79.1±1.5 81.8±0.7 73.9±1.2 88.3±0.4 78.8±1.8
AFL TWAFL 65.7±1.0 48.1±2.3 71.4±0.4 70.1±1.2 78.4±1.4 67.3±0.8 83.5±0.7 69.2±3.6
FedBuff 75.1±0.4 70.4±0.4 83.9±0.3 85.9±0.4 76.5±3.6 66.2±1.3 88.2±0.3 78.5±2.8
FedSA-GCL 81.7±0.3 71.7±1.7 83.6±0.3 89.2±0.3 85.5±0.9 80.3±4.2 94.0±0.1 84.4±2.2
4.3 Method Interpretability
To answer Q2, we examine the implementation details of our method and conduct ablation studies on two representative datasets, Cora and PubMed, both of which are split using the Metis community detection algorithm. The objective of these experiments is to validate the following points: Q2-1.The locally smoothed confidence (LSC) aggregation weights, which incorporate staleness-aware computation, can accelerate convergence in asynchronous federated graph learning. Q2-2.The proposed ClusterCast mechanism is better suited for asynchronous federated graph scenarios and improves overall accuracy. Q2-3.Our proposed clustering algorithm based on the SFM (Soft Label Feature Matrix) can effectively group clients and improve overall convergence speed. To provide a more intuitive comparison between our SFM-based clustering algorithm and the clustering algorithm in FedGTA, we visualize similarity heatmaps between clients at different client trips (100, 300, 500, and 1900). All results are shown in Fig. LABEL:fig:ablation_study and Fig. LABEL:fig:ablation_study2.
Since model evaluation is performed locally on each client’s own data, it is normal to observe a phenomenon where the accuracy initially drops before converging. This is because, at the beginning of training, the model is optimized for its local data and thus achieves high accuracy. However, after aggregating with other clients’ models, the specificity of local data decreases, leading to a temporary drop in performance. As training progresses, the model gradually converges and eventually surpasses its initial accuracy.
Regarding Q2-1, this phenomenon is demonstrated by the orange and purple curves in Fig. LABEL:fig:Training_Curve_Cora and Fig. LABEL:fig:Training_Curve_PubMed. Without the incorporation of staleness-aware locally smoothed confidence (LSC) aggregation weights, the influence of outdated models leads to unstable convergence in the later training stages, which slows down the overall convergence process.
For Q2-2, the advantage of the ClusterCast mechanism is illustrated by comparing the orange and green curves in Fig. LABEL:fig:Training_Curve_Cora and Fig. LABEL:fig:Training_Curve_PubMed. The results clearly show that incorporating ClusterCast significantly improves both the convergence speed and final accuracy of the model.
As for Q2-3, this can be observed through the orange and red curves in Fig. LABEL:fig:Training_Curve_Cora and Fig. LABEL:fig:Training_Curve_PubMed, as well as from Fig. LABEL:fig:Similarity_Cora and Fig. LABEL:fig:Similarity_PubMed. By comparing client similarity heatmaps across different client trips, it can be seen that the clustering behavior of FedGTA remains largely consistent across client trips, while our SFM-based clustering method responds more dynamically. Notably, in the PubMed dataset, FedGTA fails to form effective client clusters. Although FedGTA’s clustering appears more confident in the early training stages—resulting in faster initial convergence due to more accurate client similarity estimation—our method gradually improves its clustering confidence over time. As a result, model accuracy stabilizes and ultimately converges. This is especially evident in the Fig. LABEL:fig:Training_Curve_Cora, which show that our clustering method achieves slightly higher accuracy on Cora, and comparable accuracy on PubMed.In Fig LABEL:fig:Training_Curve_PubMed, our method underperforms FedGTA in a certain segment, which can be attributed to the inherent limitation of clustering-based learning—namely, the possible loss of valuable information. However, as shown in Fig LABEL:fig:Similarity_PubMed, FedGTA doesn’t encounter this issue, possibly because its clustering fails to produce meaningful groupings.
4.4 Robustness Comparison
To answer Q3, we conduct two groups of robustness experiments on all seven graph datasets using Metis to split client subgraphs. Specifically, we introduce two types of perturbations: random label sparsity and random topology sparsity. Each baseline is trained five times, and we report the mean performance along with the 95% confidence interval. Under the random topology sparsity setting, our method outperforms the best-performing baseline by an average of 3.1%. In the random label sparsity scenario, the improvement reaches 4.0%. These results further demonstrate the robustness of our proposed method against both structural and annotation perturbations. Detailed results are summarized in Tables 4 and 5, where the best results are highlighted using bold.
Table 4: Performance under Random Label Sparsity
type Method Cora CiteSeer PubMed CS Photo Computers Physics Overall
FedAvg 63.6±2.2 70.6±0.4 81.2±1.3 72.4±4.8 79.6±3.0 59.9±3.5 87.5±0.0 73.5±3.3
FL FedProx 61.0±3.4 70.2±0.6 81.9±0.5 74.0±3.7 76.6±6.3 55.3±2.9 87.9±0.6 72.4±3.8
MOON 64.0±5.9 64.5±7.0 70.6±2.1 65.5±3.4 72.4±6.5 59.4±3.8 87.9±0.5 69.2±3.2
FedSage+ 73.9±1.9 69.9±0.8 82.1±0.8 78.8±1.2 80.8±0.3 66.8±7.0 89.5±2.9 77.4±2.6
FGL FedGTA 71.0±1.7 71.0±0.1 77.5±2.1 89.4±0.3 82.6±0.5 73.0±1.1 94.0±0.2 79.8±3.0
FedTAD 55.1±9.2 72.3±0.5 70.6±6.3 72.9±5.6 72.9±7.6 63.0±0.8 87.1±0.6 70.6±3.5
FedAsyn 77.5±1.1 69.4±0.4 74.8±1.0 78.6±0.8 81.9±0.7 74.4±1.6 88.3±0.5 77.8±2.0
AFL TWAFL 65.5±0.7 46.0±1.1 71.4±0.2 70.0±1.9 78.8±1.4 67.9±3.0 82.8±0.4 68.9±3.8
FedBuff 75.6±0.3 69.4±0.2 83.8±0.6 85.6±1.2 79.3±1.4 67.5±1.8 88.5±1.0 78.5±2.6
FedSA-GCL 82.1±0.4 70.2±0.4 83.9±0.9 88.8±0.4 84.9±0.7 82.6±2.0 94.2±0.1 83.8±2.4
Table 5: Performance under Random Topology Sparsity
type Method Cora CiteSeer PubMed CS Photo Computers Physics Overall
FedAvg 65.0±4.1 71.3±0.3 82.2±0.6 72.3±5.9 75.4±4.4 59.8±5.5 88.0±0.5 73.4±3.2
FL FedProx 60.0±5.6 71.3±0.3 82.2±0.5 73.0±3.3 74.0±1.9 55.3±3.0 87.9±0.5 72.0±3.8
MOON 62.0±1.4 71.6±0.7 71.0±1.7 67.3±0.8 74.6±7.3 58.9±2.8 87.3±0.3 70.4±3.1
FedSage+ 74.8±1.7 69.9±0.6 81.8±1.7 79.0±0.7 79.1±2.0 69.8±2.5 88.3±0.3 77.5±2.2
FGL FedGTA 73.5±2.5 71.4±0.3 80.9±1.2 89.5±0.2 83.6±1.6 72.8±1.2 94.1±0.1 80.8±2.9
FedTAD 52.0±8.6 71.7±0.5 71.6±7.0 68.4±2.5 73.4±8.4 61.4±4.3 88.4±3.1 69.5±3.9
FedAsyn 78.2±0.5 71.0±0.3 76.5±0.8 78.4±0.6 81.8±1.2 74.4±0.7 88.4±0.3 78.4±1.8
AFL TWAFL 64.5±1.5 47.1±1.8 71.5±0.4 69.3±0.7 78.7±1.7 66.3±2.6 83.3±0.8 68.7±3.8
FedBuff 75.5±0.5 70.6±0.3 84.0±0.3 84.8±2.0 77.4±3.4 67.3±1.8 88.2±0.5 78.3±2.5
FedSA-GCL 82.2±0.4 71.0±0.4 83.9±0.6 89.0±0.2 85.2±1.0 82.0±3.3 94.0±0.2 83.9±2.3
4.5 Efficiency Analysis
To answer Q4, we follow the efficiency evaluation approach used in the semi-asynchronous algorithm FedBuff(Nguyen et al., 2022) and design an experiment to compare training efficiency. We adopt client trips as the evaluation metric, which represents the number of client communications required to reach a target validation accuracy. This metric reflects both communication and computation costs, serving as an effective proxy for training time.
The experiment is conducted on multiple graph datasets. We compare our method with strong baselines, including FedGTA (a competitive method in federated graph learning), FedBuff (a representative method in asynchronous federated learning), as well as the classical FL methods FedAvg and FedProx. For each method, we record the number of client trips required to reach the target accuracy on each dataset. All results are averaged over five runs, and the 95% confidence intervals are reported. The detailed results are shown in Tables 6 and Tables 7, where the values in parentheses indicate the number of client trips needed to achieve the specified accuracy, The best results are highlighted using bold. To provide a more intuitive comparison, we further visualize the results using 3D plots in Fig. LABEL:fig:3D_Louvain and Fig. LABEL:fig:3D_Metis.
Table 6: Client Trips Required to Reach Target Accuracy (Louvain)
Method Cora(0.64) PubMed(0.77) CS(0.77) Physics(0.88) Overall
FedAvg 1564.0±216.1 1112.0±129.8 1735.0±311.1 1920.0±0.0 1482.7±177.9
FedProx 1776.0±134.2 1164.0±124.7 1855.0±175.0 1673.3±648.4 1596.5±164.7
FedGTA 940.0±416.7 1032.0±118.6 208.0±13.6 405.0±70.6 658.9±192.1
FedBuff 556.0±41.0 533.0±70.3 1025.0±39.5 1113.0±408.5 806.8±146.1
FedSA-GCL 113.0±34.5 308.0±35.6 28.0±9.4 146.0±30.9 148.8±49.8
Table 7: Client Trips Required to Reach Target Accuracy (Metis)
Method Cora(0.56) PubMed(0.75) CS(0.67) Physics(0.87) Overall
FedAvg 1900.0±0.0 1072.0±109.1 1650.0±401.3 940.0±46.5 1237.3±205.4
FedProx 1920.0±508.2 1052.0±77.7 1533.3±366.2 1004.0±139.8 1248.0±195.8
FedGTA 428.0±59.8 1668.0±236.6 0.0±0.0 300.0±68.7 614.7±324.7
FedBuff 155.0±180.8 316.0±92.3 509.0±361.7 94.0±5.2 268.5±105.4
FedSA-GCL 69.0±23.4 452.0±54.9 0.0±0.0 0.0±0.0 130.2±90.8
The results clearly demonstrate that our proposed method consistently achieves superior efficiency across different datasets and baselines, reaching the same performance with significantly lower communication and computation overhead.
5 Conclusion
In this paper, we proposed FedSA-GCL, a novel semi-asynchronous federated graph learning framework designed to address the unique challenges posed by non-IID and graph-structured data under asynchronous settings. Our method incorporates a staleness-aware local smoothness confidence (LSC) aggregation strategy, which effectively mitigates the adverse impact of outdated models and accelerates convergence. Additionally, we introduce ClusterCast, a cluster-level broadcasting mechanism that enables inactive but structurally similar clients to benefit from timely model updates, enhancing cross-client consistency. Furthermore, we design a clustering algorithm based on the Soft Label Feature Matrix (SFM), which captures semantic similarities and demonstrates superior adaptability across training rounds.
Extensive experiments on seven real-world graph datasets, including ablation studies, robustness tests, and training efficiency evaluations, confirm that FedSA-GCL consistently outperforms state-of-the-art baselines in terms of accuracy, robustness, and convergence speed. Particularly, our method demonstrates strong resilience to label and topology sparsity, and significantly reduces communication and computation overhead measured by client trips. These results validate the effectiveness and practicality of FedSA-GCL in real-world federated graph learning scenarios.
CRediT authorship contribution statement
Zhongzheng Yuan: Conceptualization, Methodology, Software, Data Curation, Investigation, Formal analysis, Writing – original draft. Lianshuai Guo: Software, Writing – review & editing. Xunkai Li: Methodology, Visualization, Supervision, Writing – review & editing.Yinlin Zhu: Writing – review & editing. Wenyu Wang: Supervision. Meixia Qu: Supervision.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
This work was supported by the Shenzhen Fundamental Research Program, China under Grant JCYJ20230807094104009.
Data availability
The data used in this research are publicly available graph datasets.

### v2

@@
 To address these challenges, we propose FedSA-GCL, a semi-asynchronous federated framework that leverages both inter-client label distribution divergence and graph topological characteristics through a novel ClusterCast mechanism for efficient training.
-We evaluate FedSA-GCL on multiple real-world graph datasets using the Louvain and Metis split algorithm, and compare it against 9 baselines.
+We evaluate FedSA-GCL on multiple real-world graph datasets using the Louvain and Metis algorithms, and compare it against 9 baselines.
 Extensive experiments demonstrate that our method achieves strong robustness and outstanding efficiency, outperforming the baselines by an average of 2.92% with the Louvain and by 3.4% with the Metis.
@@
 Federated graph learning , Asynchronous federated learning , Non-IID data , Client clustering
-journal: Expert Systems with Applications
+journal:
 […]
@@
 Achieves +3.4% accuracy and +3.1% robustness over baseline.
-Achieves 484.5 fewer trips over baseline, improving communication efficiency.
+Reduces client trips by 484.5 over baseline, improving communication efficiency.
 1 Introduction
-Graphs are data structure for modeling complex relational systems across a wide range of domains, including social networks(Guo & Wang, 2020), drug discovery(Liu et al., 2022)(Wang et al., 2023), and recommendation platforms(He et al., 2023)(Yu et al., 2022).
+Graphs are data structure for modeling complex relational systems across a wide range of domains, including social networks[1], drug discovery[2, 3], and recommendation platforms[4, 5].
 Although graph machine learning has made significant progress in recent years, most existing methods rely on centralized storage of large-scale graph data on a single server.
 However, with growing concerns about data security and user privacy, this centralized model is often impractical in real-world scenarios.
-In practice, graph data is typically distributed across multiple data owners, and cross-site data collection becomes impractical due to privacy restrictions(Fu et al., 2022).
+In practice, graph data is typically distributed across multiple data owners, and cross-site data collection becomes impractical due to privacy restrictions[6].
 To address this challenge, the concept of Federated Graph Learning (FGL) has emerged, which integrates federated learning (FL) with graph learning.
 FGL enables collaborative model training among distributed graph data holders without requiring raw data exchange, thereby preserving privacy.
-In FGL, graph data is typically partitioned into local subgraphs held by different clients, such as transportation subnet works (Zhang et al., 2021b; Li & Zhu, 2021), social communities (Guo & Wang, 2020), or financial chains (Yang et al., 2021).
-A key challenge for FGL is subgraph heterogeneity, which refers to the significant divergence in label distributions and graph structures—driven by strong node dependencies—across clients(Baek et al., 2023; Xie et al., 2021), thereby undermining cross-client model generalization.
+In FGL, graph data is typically partitioned into local subgraphs held by different clients, such as transportation subnet works [7, 8], social communities [1], or financial chains [9].
+A key challenge for FGL is subgraph heterogeneity, which refers to the significant divergence in label distributions and graph structures—driven by strong node dependencies—across clients[10, 11], thereby undermining cross-client model generalization.
 To mitigate the impact of subgraph heterogeneity, recent studies in FGL have begun to explicitly account for the unique structural and distributional characteristics of graph data.
-For example, several works incorporate topology-aware designs to capture local graph semantics (Zhang et al., 2021a; Li et al., 2024b), while others explore personalized modeling or representation-level adaptation to better handle client-specific variations (Tan et al., 2022; Baek et al., 2023; Li et al., 2024a).
-Despite their effectiveness, these methods are still primarily designed under synchronous training protocols, which pose significant bottlenecks to efficiency in practical settings (Jiang et al., 2022).
+For example, several works incorporate topology-aware designs to capture local graph semantics [12, 13], while others explore personalized modeling or representation-level adaptation to better handle client-specific variations [14, 10, 15].
+Despite their effectiveness, these methods are still primarily designed under synchronous training protocols, which pose significant bottlenecks to efficiency in practical settings [16].
 Consequently, current FGL research continues to face two major limitations.
 L1: Synchronous FGL Suffers from Inefficiency.
-Although existing FGL methods have demonstrated promising results (Baek et al., 2023; Li et al., 2024b, a; Zhu et al., 2024), they are predominantly designed for idealized synchronous settings.
+Although existing FGL methods have demonstrated promising results [10, 13, 15, 17], they are predominantly designed for idealized synchronous settings.
 However, their reliance on synchronization raises serious concerns for practical deployment.
-In heterogeneous environments, such as those involving edge devices or clients with limited computing capacity, stragglers can delay the entire training process, leading to prolonged communication rounds and underutilized system resources (Chen et al., 2019).
+In heterogeneous environments, such as those involving edge devices or clients with limited computing capacity, stragglers can delay the entire training process, leading to prolonged communication rounds and underutilized system resources [18].
 Additionally, in real-world scenarios, client participation is often sporadic and inconsistent.
@@
 L2: Asynchronous Settings Amplify Graph Non-IID.
-Existing asynchronous federated learning (AFL) methods (Xie et al., 2019; Xu et al., 2023; Ma et al., 2021; Nguyen et al., 2022) are primarily developed for domains such as computer vision and natural language processing.
+Existing asynchronous federated learning (AFL) methods [19, 20, 21, 22] are primarily developed for domains such as computer vision and natural language processing.
 These approaches typically assume independent and identically distributed (IID) data, and lack dedicated mechanisms to address the cross-node dependencies and topological heterogeneity intrinsic to graph data.
-Although SWESALT (Liao et al., 2023) introduces asynchronous communication into FGL, it does not explicitly address the Non-IID nature of graph-structured data.
+Although SWESALT [23] introduces asynchronous communication into FGL, it does not explicitly address the Non-IID nature of graph-structured data.
 The inherent non-IID nature of graph data arises from topological heterogeneity and cross-node dependencies.
@@
 2.1 Problem Formulation
-Due to the lack of publicly available benchmark datasets for federated graph learning (FGL), existing works (Zhang et al., 2021a; Chen et al., 2024; Baek et al., 2023) typically simulate FGL scenarios by splitting a single global graph into multiple subgraphs.
+Due to the lack of publicly available benchmark datasets for federated graph learning (FGL), existing works [12, 24, 10] typically simulate FGL scenarios by splitting a single global graph into multiple subgraphs.
 Given a global graph $G=(\mathcal{V},\mathcal{E})$ with $|\mathcal{V}|=n$ nodes and $|\mathcal{E}|=m$ edges, its structure can be represented by an adjacency matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$.
 Each node $v\in\mathcal{V}$ is associated with a feature vector $\mathbf{x}_{v}\in\mathbb{R}^{f}$, forming the feature matrix $\mathbf{X}\in\mathbb{R}^{n\times f}$, and a one-hot label vector $\mathbf{y}_{v}\in\mathbb{R}^{|\mathcal{Y}|}$, resulting in the label matrix $\mathbf{Y}\in\mathbb{R}^{n\times|\mathcal{Y}|}$, where $|\mathcal{Y}|$ is the number of classes.
-To simulate a federated environment, community detection algorithms such as Metis(Karypis & Kumar, 1998) or Louvain(Blondel et al., 2008) are applied to split the global graph $G$ into $N$ subgraphs.
+To simulate a federated environment, community detection algorithms such as Metis[25] or Louvain[26] are applied to split the global graph $G$ into $N$ subgraphs.
 Each client $i\in\left\{1,\dots,N\right\}$ is assigned a local subgraph $G_{i}=(\mathcal{V}_{i},\mathcal{E}_{i})$, along with the corresponding local adjacency matrix $\mathbf{A}_{i}\in\mathbb{R}^{|\mathcal{V}_{i}|\times|\mathcal{V}_{i}|}$, feature matrix $\mathbf{X}_{i}\in\mathbb{R}^{|\mathcal{V}_{i}|\times f}$, and label matrix $\mathbf{Y}_{i}\in\mathbb{R}^{|\mathcal{V}_{i}|\times|\mathcal{Y}|}$.
 2.2 Federated Graph Learning
-FedAvg (McMahan et al., 2017) is the most widely used optimization strategy in federated learning.
+FedAvg [27] is the most widely used optimization strategy in federated learning.
 It performs global model aggregation by computing a simple weighted average of the model parameters received from participating clients, where the weights are typically proportional to the size of local training data.
-In this setting, we integrate graph neural networks (GNNs)(Kipf & Welling, 2016) with FedAvg.
+In this setting, we integrate graph neural networks (GNNs)[28] with FedAvg.
 The aggregation process at round $t$ is given by Eq.
@@
 Although FedAvg is widely adopted, it performs poorly under non-IID client data and fails to capture the unique properties of graph-structured data.
-To address this, FedSage+(Zhang et al., 2021a) integrates GraphSAGE into the federated setting and reconstructs missing inter-subgraph edges, but mainly focuses on structural completion.
-FedProto(Tan et al., 2022) tackles label distribution heterogeneity via prototype sharing, yet ignores graph topology.
-FedPub (Baek et al., 2023) introduces functional embeddings and sparse aggregation to enable personalized training on heterogeneous subgraphs.
-FedGTA (Li et al., 2024b) and AdaFGL (Li et al., 2024a) incorporate topological confidence and structure-aware personalization strategies, improving both convergence and adaptability.
-More recently, FedTAD (Zhu et al., 2024) adopts topology-aware, data-free distillation to enable structure-sensitive aggregation without data sharing.
+To address this, FedSage+[12] integrates GraphSAGE into the federated setting and reconstructs missing inter-subgraph edges, but mainly focuses on structural completion.
+FedProto[14] tackles label distribution heterogeneity via prototype sharing, yet ignores graph topology.
+FedPub [10] introduces functional embeddings and sparse aggregation to enable personalized training on heterogeneous subgraphs.
+FedGTA [13] and AdaFGL [15] incorporate topological confidence and structure-aware personalization strategies, improving both convergence and adaptability.
+More recently, FedTAD [17] adopts topology-aware, data-free distillation to enable structure-sensitive aggregation without data sharing.
 Despite these advancements, most federated graph learning (FGL) methods rely on synchronous communication, where global aggregation is blocked until all selected clients complete local training.
@@
 2.3 Asynchronous Federated Learning
-To overcome the inefficiencies of synchronous federated learning under system heterogeneity, Asynchronous Federated Learning (AFL) has been proposed to enable non-blocking client-server communication (Xu et al., 2023).
+To overcome the inefficiencies of synchronous federated learning under system heterogeneity, Asynchronous Federated Learning (AFL) has been proposed to enable non-blocking client-server communication [20].
 AFL can be broadly categorized based on its aggregation strategy into fully asynchronous and semi-asynchronous approaches.
-Fully asynchronous methods such as FedAsync (Xie et al., 2019) and TWAFL (Chen et al., 2019) allow the server to update the global model immediately upon receiving a single client update, improving reactivity but risking instability under non-IID conditions.
-In contrast, semi-asynchronous methods like FedSA (Ma et al., 2021) and FedBuff (Nguyen et al., 2022) trigger aggregation only after collecting a certain number of updates, striking a balance between convergence stability and communication efficiency.
-TimelyFL (Zhang et al., 2023) introduces a tier-based scheduling mechanism that adapts local training based on device availability and latency, further enhancing responsiveness in heterogeneous environments.
-Although some preliminary efforts, such as SWESALT (Liao et al., 2023), have extended semi-asynchronous strategies to graph data, they do not incorporate graph-specific optimizations and overlook the non-IID nature of graph data.
+Fully asynchronous methods such as FedAsync [19] and TWAFL [18] allow the server to update the global model immediately upon receiving a single client update, improving reactivity but risking instability under non-IID conditions.
+In contrast, semi-asynchronous methods like FedSA [21] and FedBuff [22] trigger aggregation only after collecting a certain number of updates, striking a balance between convergence stability and communication efficiency.
+TimelyFL [29] introduces a tier-based scheduling mechanism that adapts local training based on device availability and latency, further enhancing responsiveness in heterogeneous environments.
+Although some preliminary efforts, such as SWESALT [23], have extended semi-asynchronous strategies to graph data, they do not incorporate graph-specific optimizations and overlook the non-IID nature of graph data.
 Current AFL methods typically assume IID input and lack mechanisms to address graph-specific challenges such as topological heterogeneity and cross-node dependencies.
@@
 $\hat{Y}=Softmax(f(\tilde{\omega_{i}};G_{i}(\mathcal{V}_{i},\mathcal{E}_{i});))$ (3)
-where $G_{i}(\mathcal{V}_{i},\mathcal{E}_{i})$ denotes the local subgraph of client i, and $f(\tilde{\omega_{i}};\cdot)$is the predictive model parameterized by the aggregated weights $\tilde{\omega}_{i}$.
+[…]
 Since connected nodes exhibit similarity in both feature distributions and labels, while different clients present distinct label distribution patterns, we construct a Soft label Feature Matrix for adjacent nodes by multiplying their soft label features based on each client’s graph data.
@@
 3.2 Model Aggregation Based on Local Smoothness Confidence and Staleness
-To quantify the topological-attribute correlation between nodes, we introduce k-step Non-parametric Label Propagation (Non-param LP), which models the relationship between target nodes and their k-hop neighbors through iterative updates(Gasteiger et al., 2018).To capture higher-order structural information, the $k$-step label propagation is given in Eq.
+To quantify the topological-attribute correlation between nodes, we introduce k-step Non-parametric Label Propagation (Non-param LP), which models the relationship between target nodes and their k-hop neighbors through iterative updates[30].To capture higher-order structural information, the $k$-step label propagation is given in Eq.
 (6).
@@
 where $\hat{Y}_{i}^{(0)}$ is the initial soft label of node $i$, $\mathcal{N}(i)$ denotes the 1-hop neighbors of node $i$, $D_{i}$ is the degree of node $i$, and $\lambda\in[0,1]$ controls the propagation balance.
-The entropy of node soft labels quantifies the prediction consistency of local models, while the local smoothing confidence characterizes their consistency with neighborhood predictions(Li et al., 2024b).
+The entropy of node soft labels quantifies the prediction consistency of local models, while the local smoothing confidence characterizes their consistency with neighborhood predictions[13].
 Specifically: Lower entropy indicates stronger prediction consistency within local graph structures, yielding more reliable learned topological features.
@@
 where $|\mathcal{V}|$ is the number of nodes, $|\mathcal{Y}|$ is the soft label dimensionality (number of classes), and $D_{i}$ is the degree of node $i$.
-In asynchronous federated learning frameworks, clients conduct local training based on different versions of the global model, resulting in model version staleness(Xie et al., 2019).
+In asynchronous federated learning frameworks, clients conduct local training based on different versions of the global model, resulting in model version staleness[19].
 To address this issue, we employ a staleness metric to quantify version discrepancy and accordingly adjust client model aggregation weights.
@@
 A key challenge in semi-asynchronous federated learning is maintaining a timely and coherent information flow under partial client participation.
-Existing methods like FedBuff(Nguyen et al., 2022) trigger aggregation upon receiving a fixed number of client updates, but only send the aggregated model to those clients who participated in that current round.
+Existing methods like FedBuff[22] trigger aggregation upon receiving a fixed number of client updates, but only send the aggregated model to those clients who participated in that current round.
 As illustrated in Fig.
@@
 Algorithm 1 FedSA-GCL: Server Procedure
-1:Initialize: $t\leftarrow 0$, $\mathcal{U}\leftarrow\emptyset$, $\mathcal{B}\leftarrow\emptyset$triangleright Global round, uploaded client set, broadcast set
+1:Initialize: $t\leftarrow 0$, $\mathcal{U}\leftarrow\emptyset$, $\mathcal{B}\leftarrow\emptyset$ $\triangleright$ Global round, uploaded client set, broadcast set
 […]
-[…]
-4:     $t\leftarrow t+1$,$\mathcal{U}\leftarrow\emptyset$$\mathcal{B}\leftarrow\emptyset$
-[…]
+3:while $\text{UploadQueue.qsize()}\geq K$ do
+4:  $t\leftarrow t+1$$\mathcal{U}\leftarrow\emptyset$$\mathcal{B}\leftarrow\emptyset$
+5:  while $\text{UploadQueue.qsize()}\neq 0$ do
 6:   Receive $(LSC_{i},SFM_{i},\omega_{i},\tau_{i},i)$ from UploadQueue.get()
@@
 Algorithm 2 FedSA-GCL: Client Procedure
+1:Initialize: $\tau\leftarrow 0$, $\texttt{DownloadQueue}\leftarrow\texttt{Queue()}$,update $\leftarrow$ null $\triangleright$ Python thread-safe queue for receiving server updates
+2:while $\texttt{DownloadQueue.qsize()}\neq 0$ do
 […]
 […]
-3:     Receive $(\tilde{\omega},\tau,\tilde{LSC})$ from DownloadQueue.get()
-4:     Replace stored update with $(\tilde{\omega},\tau,\tilde{LSC})$triangleright Keep only the latest update
 5:end while
@@
 12:Compute soft label via Eq. (3)
-13:Compute $SFM$ via Eq. (4)
+13:Compute SFMSFM via Eq. (4)
 14:Compute non-parametric LP via Eq. (6)
-15:Compute $LSC$ via Eq. (7)
+15:Compute LSCLSC via Eq. (7)
 16:Send $(\omega^{\prime},\tau,SFM,LSC,\text{client ID})$ to UploadQueue
@@
 To comprehensively evaluate the performance of our method against other baselines, we select seven graph datasets.
-Our experiments are conducted on three small-scale citation networks (Cora, CiteSeer, PubMed) (Yang et al., 2016), two medium-scale user-item datasets (Amazon Computers, Amazon Photo), and two medium-scale co-authorship networks (Coauthor CS, Coauthor Physics) (Shchur et al., 2018).
+Our experiments are conducted on three small-scale citation networks (Cora, CiteSeer, PubMed) [31], two medium-scale user-item datasets (Amazon Computers, Amazon Photo), and two medium-scale co-authorship networks (Coauthor CS, Coauthor Physics) [32].
 Detailed descriptions of these datasets and the proportions of the train/validation/test splits are presented in Table 1.
-To simulate realistic federated graph learning scenarios, we apply two community detection algorithms, Metis(Karypis & Kumar, 1998) and Louvain(Blondel et al., 2008), to split each dataset into 20 clients.
+To simulate realistic federated graph learning scenarios, we apply two community detection algorithms, Metis[25] and Louvain[26], to split each dataset into 20 clients.
 These splits are used for federated graph training.
@@
 We compare our method against nine baselines.
-For traditional federated learning (FL), we select FedAvg(McMahan et al., 2017), FedProx(Li et al., 2020), and MOON(Li et al., 2021).
-For federated graph learning (FGL), we include FedSage+(Zhang et al., 2021a), FedGTA(Li et al., 2024b), and FedTAD(Zhu et al., 2024).
-For asynchronous federated learning (AFL), we compare with FedAnsy(Xie et al., 2019), TWAFL(Chen et al., 2019), and FedBuff(Nguyen et al., 2022).
+For traditional federated learning (FL), we select FedAvg[27], FedProx[33], and MOON[34].
+For federated graph learning (FGL), we include FedSage+[12], FedGTA[13], and FedTAD[17].
+For asynchronous federated learning (AFL), we compare with FedAnsy[19], TWAFL[18], and FedBuff[22].
 All experiments are conducted using the GCN model as the backbone architecture.
 Hyperparameters.
-Our asynchronous training scripts are developed based on the open-source federated graph learning framework OpenFGL(Li et al., 2024c).
+Our asynchronous training scripts are developed based on the open-source federated graph learning framework OpenFGL[35].
 For baseline methods, we adopt the hyperparameter settings reported in their original papers.
-In federated learning, the number of communication rounds to reach a target accuracy is a common evaluation metric(Nguyen et al., 2022).
+In federated learning, the number of communication rounds to reach a target accuracy is a common evaluation metric[22].
 However, since asynchronous and synchronous approaches differ in their notion of a round, FedBuff proposes the client trip — a cycle of model download, local training for one epoch, and model upload—as a fair basis for comparison.
@@
 4.5 Efficiency Analysis
-To answer Q4, we follow the efficiency evaluation approach used in the semi-asynchronous algorithm FedBuff(Nguyen et al., 2022) and design an experiment to compare training efficiency.
+To answer Q4, we follow the efficiency evaluation approach used in the semi-asynchronous algorithm FedBuff[22] and design an experiment to compare training efficiency.
 We adopt client trips as the evaluation metric, which represents the number of client communications required to reach a target validation accuracy.